OpenStudy (anonymous):

NEED HELP PLEASE 1.) The function Tan θ = y has θ restricted to 2.) Find θ when Cos-1 0 = θ. Give your answer in degrees 3.) The domain of function Tan-1 a = θ is restricted to

OpenStudy (solomonzelman):

Tan 90 is undefined (for example)

OpenStudy (anonymous):

so would it be {θ | -90º ≤θ ≤90º}?

zepdrix (zepdrix):

Well, it's undefined `at` -90 and 90. So we need a strict inequality, can't have the `or equal to`. But yes, that gives us one interval of tangent. There are a few more +_+ infinitely many more. Umm lemme think of how we would write this..

OpenStudy (anonymous):

ohh i see

zepdrix (zepdrix):

So you're trying to write it in set notation? This is kinda sloppy.... but I dunno maybe it will help. \[\Large\rm \{\theta|-\infty\lt\theta\lt\infty,\qquad\theta\ne \frac{\pi}{2}+k\pi,\qquad k\in\mathbb Z\}\]

zepdrix (zepdrix):

So I'm saying that theta can take on EVERY value, except pi/2 and any angle that's pi/2 plus/minus a multiple of pi.

zepdrix (zepdrix):

So if k=-1. It gives us theta `cannot equal` -pi/2, yes?

zepdrix (zepdrix):

Oh you were doing this in degrees! Doh! I probably made this more difficult for you :c

OpenStudy (anonymous):

yeah haha it says to do it in degrees that's why i was processing what you were trying to say because I have not done it like that before

zepdrix (zepdrix):

\[\Large\rm \{\theta|-\infty\lt\theta\lt\infty,\qquad\theta\ne 90^{\circ}+k\cdot180^{\circ},\qquad k\in\mathbb Z\}\]Maybe it's still a little confusing :) But here's what it looks like in degrees.

zepdrix (zepdrix):